# Hidden Population III Descendants in Ultra-Faint Dwarf Galaxies

### v1

@@ Abstract
The elusive properties of the first (Pop III) stars can be indirectly unveiled by uncovering their true descendants. To this aim, we exploit our data-calibrated model for the best-studied ultra-faint dwarf (UFD) galaxy, Boötes I, which tracks the chemical evolution (from carbon to zinc) of individual stars from their formation to the present day. We explore the chemical imprint of Pop III supernovae (SNe), with different explosion energies and masses, showing that they leave distinct chemical signatures in their descendants. We find that UFDs are strongly affected by SNe-driven feedback resulting in a very low fraction of metals retained by their gravitational potential well ($<2.5\%$). Furthermore, the higher the Pop III SN explosion energy, the lower the fraction of metals retained. Thus, the probability to find descendants of energetic Pair Instability SNe is extremely low in these systems. Conversely, UFDs are ideal cosmic laboratories to identify the fingerprints of less massive and energetic Pop III SNe through their [X/Fe] abundance ratios. Digging into the literature data of Boötes I, we uncover three hidden Pop III descendants: one mono-enriched and two multi-enriched. These stars show the chemical signature of Pop III SNe in the mass range $[20-60]\>\rm{M_{\odot}}$, spanning a wide range in explosion energies $[0.3-5]10^{51}$ erg. In conclusion, Pop III descendants are hidden in ancient UFDs but those mono-enriched by a single Pop III SN are extremely rare. Thus, self-consistent models such as the one presented here are required to uncover these precious fossils and probe the properties of the first Pop III supernovae.
@@ 1 Introduction
Despite these discoveries at extremely low-metallicities, models predict that the probability to find direct descendants of high-energy Pop III SNe increases towards higher [Fe/H] (Vanni et al. 2023). However, their detection becomes more challenging at higher [Fe/H] due to the increasing numbers of subsequent generations of Pop II descendants there (Salvadori et al., 2015; de Bennassuti et al., 2017; Hartwig et al., 2018; Koutsouridou et al., 2023). For example, the descendants of PISN are expected to appear in a very broad metallicity range $\rm-4<[Fe/H]<-1$ with a peak at $\rm[Fe/H]\approx-1.8$ (Karlsson et al., 2013; de Bennassuti et al., 2017; Salvadori et al., 2019; Koutsouridou et al., 2024). Indeed, several candidate PISN descendants have been found at [Fe/H]$\approx-2$ (Aguado et al., 2023) including the one identified by Xing et al. (2023), whose abundance pattern is unfortunately not anymore consistent with a pure PISN enrichment (Skúladóttir et al. 2024; Thibodeaux et al. 2024).
@@ 2 Model description
Figure 1: Assumed EDFs: four cases in which the EDF is a top-hat function centered on each type of SNe ($100\%$ probability), and the last case in which the all SNe explosion energies are equiprobable ($25\%$ probability of each type)
Initial conditions: We assume Boötes I to evolve in isolation, i.e. without experiencing major merger events. This is consistent with cosmological models (Salvadori & Ferrara, 2009; Salvadori et al., 2015), which are used to set the initial conditions: the epoch of virialization, $\rm z_{vir}\approx 8.5$, and the dark-matter halo mass, $\rm M_{h}\approx 10^{7.3}\>\rm{M_{\odot}}$.
Star formation: At the virialization epoch we compute the total amount of gas [^1] We adopt $\Lambda$ cold dark matter ($\Lambda$CDM) with $h=0.669$, $\Omega_{\rm m}=0.3103$, $\Omega_{\Lambda}=0.6897$, $\Omega_{\rm b}h^{2}=0.02234$, $\sigma_{8}=0.8083$, $n=0.9671$, according to the latest Plank results (Planck Collaboration et al., 2018) as $M_{g}=\Omega_{b}/\Omega_{m}\>\rm M_{h}$, and we assume it to be pristine. The star-formation becomes possible when the gas begins to fall into the central part of the halo and cool down (see Rossi et al., 2021, for the infall rate). We evaluate the total mass of stars formed in a timescale of 1 Myr, assuming a star-formation rate regulated by the free-fall time ($\rm t_{ff}$) and the mass of cold gas accumulated via infall ($\rm M_{cold}$): $\rm\Psi=\epsilon_{\star}M_{cold}/t_{ff}$, where $\epsilon_{\star}$ is a free parameter of the model, equal for both Pop III and Pop II/I stars.
Following the critical metallicity scenario (Bromm et al., 2001; Schneider et al., 2003; Omukai et al., 2005), Pop III stars only form if the metallicity of the ISM gas222 $\rm Z_{gas}=M^{ISM}_{Z}/M_{gas}$ where $\rm M^{ISM}_{Z}$ and $\rm M_{gas}$ are the mass of metals and gas in the ISM, respectively., $\rm Z_{gas}$, is below the critical metallicity, $\rm Z_{cr}=10^{-4.5\pm 1}Z_{\odot}$, i.e. $\rm Z_{gas}\leq Z_{cr}$, while normal Pop II/I stars form when $\rm Z_{gas}>Z_{cr}$.
@@
where $m_{\star}$ is the stellar mass and $m_{ch}$ the characteristic mass of the IMF. For Pop III stars $m_{ch}=10\>\rm M_{\odot}$ and $m_{\star}=[0.8-1000]\>\rm{M_{\odot}}$ (consistent with the results of Hirano et al., 2014; Hirano & Bromm, 2017; Rossi et al., 2021; Pagnini et al., 2023); while for Pop II stars $m_{ch}=0.35\>\rm M_{\odot}$ and $m_{\star}=[0.1-100]\>\rm{M_{\odot}}$.
The model includes the incomplete sampling of the stellar IMF for both Pop III and Pop II stellar populations. Every time that the conditions for star formation are reached, a discrete number of stars is formed, with mass selected randomly from the assumed IMF. This is a key feature of the model, essential for accurately simulating the evolution of poorly star-forming UFDs in which the assumed IMF is never fully sampled throughout its entire evolution (see Rossi et al. 2021).
Chemical feedback: The chemical enrichment of the gas is followed by taking into account the mass-dependent timescales for the evolution of individual stars. In particular, the semi-analytical model tracks and follows the chemical elements (from carbon to zinc) released by SNe and AGB stars from both the $\rm Pop~{}III$  and $\rm Pop~{}II/I$  stellar populations.
  • SNe can enrich the ISM with all chemical elements from carbon to zinc. For $\rm Pop~{}III$  SNe in the mass range $m_{\rm Pop~{}III}=[10-100]\>\rm{M_{\odot}}$ we use the stellar yields from Heger & Woosley (2010), while for massive PISN we adopt the ones of Heger & Woosley (2002) in the mass range $m_{\text{PISN}}=[140-260]\>\rm{M_{\odot}}$. For $\rm Pop~{}II/I$  SNe in the mass range $[8-40]\>\rm{M_{\odot}}$ we use the updated metal yields of Limongi & Chieffi (2018), model R, without rotation and with a fixed explosion energy $\rm E_{SN}=10^{51}erg$.
@@
Energy Distribution Function (EDF) of Pop III SNe: For $(10-100)\>\rm{M_{\odot}}$ Pop III stars ending their lives as SNe, we follow Heger & Woosley (2010) and explore different possible explosion energies :
@@
We explore five different possibilities for the unknown EDF of Pop III SNe, shown in Fig.1. In the first four cases we assume extreme scenarios in which Pop III SNe are all of a single type: only faint, only cc, only high, or only hyper. This corresponds to a top-hat EDF in which the probability to form the chosen type of SNe is $100\%$. In the last case we assume an EDF spread among all different explosion energies (``````` `all types” model), where each SNe type has an equal probability of $25\%$. Thus, every time a Pop III SN progenitor forms we randomly assign to it an explosion energy.
@@
Conversely, in the mass range $[140-260]\>\rm{M_{\odot}}$, we assume that $\rm Pop~{}III$  stars evolve as PISNe with an explosion energy $\rm E_{SN}=[10,100]\times 10^{51}erg$, proportional to the stellar mass (Heger & Woosley 2002), and therefore independent from the assumed EDF for SNe with less massive progenitor stars. Note that the probability to form PISNe is regulated by both the assumed Pop III IMF and the star-formation rate, which affect the random sampling of the Pop III IMF (see Rossi et al., 2021). Assuming a fully sampled Pop III IMF we obtain that this probability is $\sim 0.2\%$.
Data Calibration: The model is data-calibrated: the free parameters ($\epsilon_{\star}$, $\epsilon_{w}$) have been fixed to reproduce the observed properties of the Boötes I UFD, i.e. the total luminosity, $L_{bol}$, the average stellar iron-abundance,$\langle\rm[Fe/H]\rangle$, and the star formation history (for details see Rossi et al. 2021). Due to the stochastic nature of the IMF sampling, the chemical enrichment history of Boötes I can vary between different runs of the code. For this reason, all the results presented here have been obtained by averaging over 1000 possible Boötes I evolutionary histories and by quantifying the scatter among them.
3 The energy distribution of Pop III SNe: impact
As a first step we investigate how the predicted properties of Boötes I change by assuming different EDFs of Pop III SNe, as defined in the previous Section (Fig. 1). While the predicted $\rm\langle[Fe/H]\rangle$, $L_{bol}$ and star formation history of Boötes I, are minimally influenced by the assumed EDF of Pop III SNe, as we shall now see, the Metallicity Distribution Function (MDF) is notably impacted. For this reason, we use a $\chi^{2}$ procedure to quantify which is the best model, among those assuming different EDFs, that reproduces the observed MDF of Boötes I, along with its global properties i.e. $\rm\langle[Fe/H]\rangle$, and $L_{bol}$.
Firstly, we consider that SNe in the mass range $[10-100]\>\rm{M_{\odot}}$ are all of a unique type. Then we relax this assumption and assume a flat EDF in which the four SNe types have the same probability ($25\%$). The impact of the different Pop III EDFs on the MDF of Boötes I is shown in Fig. 2. The $\chi^{2}$ values for the different models are reported in Table 1.

### v2

@@ Abstract
The elusive properties of the first (Pop III) stars can be indirectly unveiled by uncovering their true descendants. To this aim, we exploit our data-calibrated model for the best-studied ultra-faint dwarf (UFD) galaxy, Boötes I, which tracks the chemical evolution (from carbon to zinc) of individual stars from their formation to the present day. We explore the chemical imprint of Pop III supernovae (SNe), with different explosion energies and masses, showing that they leave distinct chemical signatures in their descendants. We find that UFDs are strongly affected by SNe-driven feedback resulting in a very low fraction of metals retained by their gravitational potential well ($<2.5\%$). Furthermore, the higher the Pop III SN explosion energy, the lower the fraction of metals retained. Thus, the probability to find descendants of energetic Pair Instability SNe is extremely low in these systems. Conversely, UFDs are ideal cosmic laboratories to identify the fingerprints of less massive and energetic Pop III SNe through their [X/Fe] abundance ratios. Digging into the literature data of Boötes I, we uncover three hidden candidates for Pop III descendants: one mono-enriched and two multi-enriched. These stars show the chemical signature of Pop III SNe in the mass range $[20-65]\>\rm{M_{\odot}}$, spanning a wide range in explosion energies $[0.3-10]10^{51}$ erg. In conclusion, candidates for Pop III descendants are hidden in ancient UFDs but those mono-enriched by a single Pop III SN are extremely rare. Thus, self-consistent models such as the one presented here are required to uncover these precious fossils and probe the properties of the first Pop III supernovae.
@@ 1 Introduction
Despite these discoveries at extremely low-metallicities, models predict that the probability to find direct descendants of high-energy Pop III SNe increases towards higher [Fe/H] (Vanni et al. 2023). However, their detection becomes more challenging at higher [Fe/H] due to the increasing numbers of subsequent generations of Pop II descendants there (Salvadori et al., 2015; de Bennassuti et al., 2017; Hartwig et al., 2018; Koutsouridou et al., 2023a). For example, the descendants of PISN are expected to appear in a very broad metallicity range $\rm-4<[Fe/H]<-1$ with a peak at $\rm[Fe/H]\approx-1.8$ (Karlsson et al., 2013; de Bennassuti et al., 2017; Salvadori et al., 2019; Koutsouridou et al., 2024). Indeed, several candidate PISN descendants have been found at [Fe/H]$\approx-2$ (Aguado et al., 2023) including the one identified by Xing et al. (2023), whose abundance pattern is unfortunately not anymore consistent with a pure PISN enrichment (Skúladóttir et al. 2024; Thibodeaux et al. 2024).
@@ 2 Model description
Figure 1: Assumed EDFs for Pop III (top panel): four cases in which the EDF is a top-hat function centered on each type of SNe ($100\%$ probability), and the last case in which the all SNe explosion energies are equiprobable ($25\%$ probability of each type). Assumed EDF for Pop II/I stars (bottom panel): the first case assumes a top-hat EDF centered on the explosion energy of core-collapse (CC) supernovae (100% probability); while the second case considers a mixed distribution where $70\%$ of Pop II/I SNe have CC energies, and $30\%$ correspond to hypernovae.
Initial conditions: We assume Boötes I to evolve in isolation, i.e. without experiencing major merger events. This is consistent with cosmological models (Salvadori & Ferrara, 2009; Salvadori et al., 2015), which are used to set the initial conditions: the epoch of virialization, $\rm z_{vir}\approx 8.5$, and the corresponding dark matter halo mass at that redshift, $\rm M_{h}\approx 10^{7.3}\>\rm{M_{\odot}}$.
Star formation: At the virialization epoch we compute the total amount of gas [^1] We adopt $\Lambda$ cold dark matter ($\Lambda$CDM) with $h=0.669$, $\Omega_{\rm m}=0.3103$, $\Omega_{\Lambda}=0.6897$, $\Omega_{\rm b}h^{2}=0.02234$, $\sigma_{8}=0.8083$, $n=0.9671$, according to the latest Plank results (Planck Collaboration et al., 2018) as $M_{g}=\Omega_{b}/\Omega_{m}\>\rm M_{h}$, and we assume it to be pristine. The star-formation becomes possible when the gas begins to fall into the central part of the halo and cool down (see Rossi et al., 2021, for the infall rate). We evaluate the total mass of stars formed in a timescale of 1 Myr, assuming a star-formation rate regulated by the redshift dependent free-fall time ($\rm t_{ff}(z)$) and the mass of cold gas accumulated via infall ($\rm M_{cold}$): $\rm\Psi=\epsilon_{\star}M_{cold}/t_{ff}(z)$, where $\epsilon_{\star}$ is a free parameter of the model, equal for both Pop III and Pop II/I stars. Note that since the free-fall time $\rm t_{\text{ff}}(z)$ is larger at high redshift (i.e., when Pop III stars form), this results in a smaller star formation rate ($\Psi$) for Pop III stars compared to Pop II/I stars, even if $\epsilon_{\star}$ is the same.
Following the critical metallicity scenario (Bromm et al., 2001; Schneider et al., 2003; Omukai et al., 2005), Pop III stars only form if the metallicity of the ISM gas222 $\rm Z_{gas}=M^{ISM}_{Z}/M_{gas}$ where $\rm M^{ISM}_{Z}$ and $\rm M_{gas}$ are the mass of metals and gas in the ISM, respectively., $\rm Z_{gas}$, is below the critical metallicity, $\rm Z_{cr}=10^{-4.5\pm 1}Z_{\odot}$, i.e. $\rm Z_{gas}\leq Z_{cr}$, while normal Pop II stars form when $\rm Z_{gas}>Z_{cr}$.
@@
where $m_{\star}$ is the stellar mass and $m_{ch}$ the characteristic mass of the IMF. For Pop III stars $m_{ch}=10\>\rm M_{\odot}$ and $m_{\star}=[0.8-1000]\>\rm{M_{\odot}}$ (consistent with the results of Hirano et al., 2014; Hirano & Bromm, 2017; Rossi et al., 2021; Pagnini et al., 2023); while for Pop II/I stars $m_{ch}=0.35\>\rm M_{\odot}$ and $m_{\star}=[0.1-100]\>\rm{M_{\odot}}$.
The model includes the incomplete sampling of the stellar IMF for both Pop III and Pop II/I stellar populations. Every time that the conditions for star formation are reached, a discrete number of stars is formed, with mass selected randomly from the assumed IMF. This is a key feature of the model, essential for accurately simulating the evolution of poorly star-forming UFDs in which the assumed IMF is never fully sampled throughout its entire evolution (see Rossi et al. 2021).
Chemical feedback: The chemical enrichment of the gas is followed by taking into account the mass-dependent timescales for the evolution of individual stars. Stellar lifetimes are calculated using the relation from Raiteri et al. (1996), which accounts for both stellar mass and metallicity, reflecting the chemical composition of the ISM from which the stars originated. The semi-analytical model tracks and follows the chemical elements (from carbon to zinc) released by SNe and AGB stars from both the $\rm Pop~{}III$  and $\rm Pop~{}II/I$  stellar populations.
  • SNe can enrich the ISM with all chemical elements from carbon to zinc. For $\rm Pop~{}III$  SNe in the mass range $m_{\rm Pop~{}III}=[10-100]\>\rm{M_{\odot}}$ we use the stellar yields from Heger & Woosley (2010), while for massive PISN we adopt the ones of Heger & Woosley (2002) in the mass range $m_{\text{PISN}}=[140-260]\>\rm{M_{\odot}}$. For $\rm Pop~{}II/I$  SNe in the mass range $[8-40]\>\rm{M_{\odot}}$ we use three set of stellar yields: Limongi & Chieffi (2018) (hereafter LC), specifically model R, which assumes no rotation and a fixed explosion energy of $\rm E_{SN}=10^{51}erg$; the yields from Nomoto et al. (2013), including only core-collapse Pop II SNe with $\rm E_{SN}=10^{51}erg$ (hereafter NK CC); and Nomoto et al. (2013) models with hypernove (hereafter NK CC+HN) with $\rm E_{SN}=[10-30]\times 10^{51},erg$ depending on the stellar mass in the range [13-40]$\>\rm{M_{\odot}}$.
@@
  • Type Ia supernovae (SNIa): The contribution of SNIa has been included by adopting the yields by Iwamoto et al. (1999) and the bimodal delay time distribution observationally derived by Mannucci et al. (2006). At each time-step and for each stellar burst we computed the rate of SNIa by following Matteucci et al. (2006), where the normalization constant has been fixed to reproduce the actual rate of SNIa in the Milky Way.
@@
Radiative feedback: Our model does not explicitly incorporate radiative feedback processes; however, the halo mass selected for the Boötes I dwarf galaxy is derived from cosmological models that account for the increase in the minimum mass required for star formation due to the influence of the Lyman-Werner background. The model do not account for internal radiative feedback from young stars which may have a role in reducing the efficiency of star formation of low-mass galaxies (Wise et al. (2012); Jeon et al. (2014)).
Energy Distribution Function (EDF) of Pop III SNe: For $[10-100]\>\rm{M_{\odot}}$ Pop III stars ending their lives as SNe, we follow Heger & Woosley (2010) and explore different possible explosion energies :
@@
We explore five different possibilities for the unknown EDF of Pop III SNe, shown in Fig.1 (top panel). In the first four cases we assume extreme scenarios in which Pop III SNe are all of a single type: only faint, only cc, only high, or only hyper. This corresponds to a top-hat EDF in which the probability to form the chosen type of SNe is $100\%$. In the fifth case we assume an EDF spread among all different explosion energies (``````` `all types” model), where each SNe type has an equal probability of $25\%$. Thus, every time a Pop III SN progenitor forms we randomly assign to it an explosion energy.
@@
Conversely, in the mass range $[140-260]\>\rm{M_{\odot}}$, we assume that $\rm Pop~{}III$  stars evolve as PISNe with an explosion energy $\rm E_{SN}=[10,100]\times 10^{51}erg$, proportional to the stellar mass (Heger & Woosley 2002), and therefore independent from the assumed EDF for SNe with less massive progenitor stars. Note that the probability to form PISNe is regulated by both the assumed Pop III IMF and the star-formation rate, which affect the random sampling of the Pop III IMF (see Rossi et al., 2021). Assuming a fully sampled Pop III IMF (with $m_{ch}=10\>\rm{M_{\odot}}$) we obtain that this probability is $\sim 0.2\%$.
EDF of Pop II/I SNe: For Pop II/I stars that end their lives as SNe, we explore the two EDFs shown in Fig.1 (bottom panel):
  • core-collapse: A flat EDF where all Pop II/I SNe have a fixed explosion energy of $\rm E_{SN}=10^{51}$ erg. The stellar yields associated to this EDF are derived from either Limongi & Chieffi (2018) or Nomoto et al. (2013). In the LC model, the stellar SNe mass range is $[8-40],\>\rm{M_{\odot}}$, while for the NK CC model, it is $[13-40],\>\rm{M_{\odot}}$.
  • core-collapse + hypernovae: We also consider an EDF where $70\%$ of Pop II/I SNe are core-collapse supernovae with a fixed explosion energy of $\rm E_{SN}=10^{51}$ erg, while the remaining $30\%$ are hypernovae Ishigaki et al. (2014). The hypernovae explosion energy is mass-dependent and spans the range $\rm E_{SN}=[10-30]\times 10^{51}$ erg.
$\rm\log(L_{\star}/L_{\odot})$ $t_{50}\rm(Myr)$ $\rm\chi^{2}_{MDF}$ $(\epsilon_{\star},\epsilon_{w})$ $\chi^{2}_{\nu}$
LC CC yields
FAINT $4.69\pm 0.03$ $24\pm 1$ 186 (1.0, 0.009) 25.92
CC $4.78\pm 0.05$ $29\pm 3$ 40 (1.0, 0.008) 12.57
HIGH $4.61\pm 0.03$ $33\pm 1$ 23 (0.5, 0.0014) 5.71
HYPER $4.58\pm 0.26$ $37\pm 2$ 37 (0.5, 0.001) 6.20
ALL TYPE $4.62\pm 0.06$ $33\pm 2$ 18 (0.9, 0.002) 5.05
NK CC yields
FAINT $5.29\pm 0.01$ $33\pm 1$ 80 (0.5, 0.002) 71.54
CC $5.32\pm 0.05$ $34\pm 1$ 21 (1.2, 0.001) 56.16
HIGH $5.27\pm 0.11$ $34\pm 1$ 30 (1.2, 0.001) 31.05
HYPER $4.34\pm 0.35$ $37\pm 2$ 89 (0.5, 0.001) 12.17
ALL TYPE $4.97\pm 0.04$ $35\pm 1$ 24 (0.4, 0.002) 23.47
NK CC+HN yields
FAINT $5.24\pm 0.02$ $31\pm 1$ 77 (1.2, 0.002) 63.32
CC $5.28\pm 0.05$ $33\pm 1$ 48 (1.1, 0.001) 56.27
HIGH $5.24\pm 0.10$ $33\pm 1$ 26 (1.2, 0.001) 33.30
HYPER $4.89\pm 0.10$ $34\pm 1$ 42 (0.5, 0.004) 14.03
ALL TYPE $5.06\pm 0.07$ $33\pm 1$ 14 (0.6, 0.001) 25.00
Table 1: Simulated properties of Boötes I with corresponding $\chi^{2}$ values, using using Pop III yields from Heger & Woosley (2010); and different Pop I/II yields as listed. The EDFs are adopted from Fig. 1). The observed properties of Boötes I include the total luminosity, $\log(L_{\star}/L_{\odot})=4.5\pm 0.1$ (Kirby et al., 2013), the time needed to form $50\%$ of the total stellar mass ($t_{50}$, Brown et al. 2014), and the MDF (Jenkins et al., 2021), see text for details.
3 Model calibration
The model has been calibrated using observational data as detailed in (Rossi et al., 2021). This means that the two free parameters ($\epsilon_{\star}$, $\epsilon_{w}$) are fixed to match the observed properties of Boötes I: the total luminosity, $\rm\log(L_{\rm obs}/L_{\odot})=4.5\pm 0.1$ (Kirby et al., 2013); the Metallicity Distribution Function (MDF) (Jenkins et al. (2021), which is known to be one of the key observables to constrain chemical evolution models (e.g., Salvadori et al., 2007; Koutsouridou et al., 2023b); and the time needed to form 50% of the total stellar mass, $t_{50}$, which provides the weakest constraints ranging from less than $100$ Myr to $\approx 1$ Gyr (Brown et al., 2014). Compared to Rossi et al. (2021), we have updated the Boötes I stellar sample using data from Jenkins et al. (2021), consisting of 54 stars. This results in a revised mean metallicity of $\langle\rm[Fe/H]\rangle=-2.34\pm 0.28$.
Due to the stochastic nature of the IMF sampling, the chemical enrichment history of Boötes I varies across different runs of the simulation. To account for this variability, all results presented here are derived by averaging over 1000 independent realizations of the galaxy’s evolutionary history. The associated scatter across these runs is quantified as the standard deviation.
From the ensemble of realizations, we compute the mean simulated values ($\langle\rm X_{\rm sim}\rangle$) and their associated uncertainties ($\sigma_{\rm sim}$) for the observables mentioned above (luminosity, MDF, and $t_{50}$). These values are used to compute the total $\chi^{2}$ as the sum of three terms, each corresponding to a specific observable:
$\chi^{2}=\chi^{2}_{\rm Lum}+\chi^{2}_{t_{50}}+\chi^{2}_{\rm MDF}.$ (2)
The terms $\chi^{2}_{\rm Lum}$ and $\chi^{2}_{t_{50}}$ are computed using the general formula:
$\chi^{2}_{\rm X}=\frac{\left(X_{\rm obs}-\langle X_{\rm sim}\rangle\right)^{2}}{\sigma_{\rm X,obs}^{2}+\sigma_{\rm X,sim}^{2}},$ (3)
where $X$ represents the observable: total luminosity (${\rm Lum}$) or $t_{50}$. The observed values ($X_{\rm obs}$) are compared to the mean simulated values ($\langle X_{\rm sim}\rangle$), and the combined uncertainties from observations ($\sigma_{\rm X,obs}$) and simulations ($\sigma_{\rm X,sim}$) are used in the denominator.
The third contribution to the total $\chi^{2}$ is derived from the MDF. For each run of the simulation, we compute the MDF by binning stars in intervals of [Fe/H]. The mean simulated MDF ($\langle{\rm MDF}{\rm sim}\rangle$) is then calculated by averaging the number of stars in each iron bin across 1000 independent realizations, and the associated uncertainty is quantified as the standard deviation among these runs. Similarly, the observational MDF is characterized by the observed number of stars in each [Fe/H] bin ($N{\rm obs}$), with the uncertainty on $N_{\rm obs}$ in each bin assumed to be $\sigma_{\rm MDF,obs}=\sqrt{N_{\rm obs}}$. The $\chi^{2}_{\rm MDF}$ term quantifies the agreement between the simulated and observed MDF and is computed as:
$\chi^{2}_{\rm MDF}=\sum_{k=1}^{n}\frac{\left(N_{\rm obs}[k]-\langle N_{\rm sim}[k]\rangle\right)^{2}}{\sigma_{\rm MDF,obs}[k]^{2}+\sigma_{\rm MDF,sim}[k]^{2}},$ (4)
where $n$ is the number of [Fe/H] bins, $N_{\rm obs}[k]$ is the number of observed stars in the $k$-th bin, $\langle N_{\rm sim}[k]\rangle$ is the mean simulated number of stars in the same bin, and $\sigma_{\rm MDF,obs}[k]$ and $\sigma_{\rm MDF,sim}[k]$ are the uncertainties associated with the observed and simulated MDF, respectively. In this work, we include all bins with $\rm[Fe/H]>-4$. Finally, we computed the reduced chi-square333 $\chi^{2}_{\nu}=\chi^{2}/\nu$, where $\nu$ is the number of degrees of freedom, calculated as the total number of data points ($n_{\rm data}$) minus the number of free parameters in the model. In this work, $\nu=n_{\rm data}-2=9$, considering the two free parameters ($\epsilon_{\star}$, $\epsilon_{w}$) in the model , $\chi^{2}_{\nu}$, to assess the goodness of fit of the model. Finally, we fix the free parameters ($\epsilon_{\star}$, $\epsilon_{w}$) by minimizing $\chi^{2}_{\nu}$.
Figure 2: Comparison of the observed (black outlines) and simulated (shaded colors) MDFs of Boötes I. Columns show different assumptions for the Pop II/I EDFs/yields, while rows show different Pop III EDFs.
4 Fiducial Model
[…]
For each combination of Pop II stellar yields and EDFs, we calibrated the model by determining the pair of parameters ($\epsilon_{\star}$, $\epsilon_{\text{w}}$) that minimizes the reduced chi-squared ( $\chi^{2}_{\nu}$ ) between theoretical predictions and observational constraints. Table 1 summarizes the results and lists the simulated values of $\rm\log(L_{\star}/L_{\odot})$, $t_{50}$, the contribution to the $\chi^{2}$ from the MDF, $\chi^{2}_{\rm MDF}$, and the $\epsilon_{\star}$, $\epsilon_{w}$ values that minimize the reduced chi-square and $\chi^{2}_{\nu}$.
Our results show that the bulk of the chemical enrichment in Boötes I is dominated by Pop II/I stars. However, Pop III stars determine the initial metallicity flow and chemical enrichment conditions on which Pop II/I stars subsequently form, and therefore have a significant impact on the overall shape of the total MDF.
Nonetheless, we observe that to achieve a good fit as the explosion energy of Pop III stars increases, it becomes necessary to either (1) reduce the star formation efficiency ($\epsilon_{\star}$), thereby forming fewer Pop III SNe, or (2) decrease the wind efficiency ($\epsilon_{\text{w}}$), allowing more chemical elements from the energetic Pop III SNe to be retained within the system. However, a clear trend is not always evident due to the stochastic nature of early galaxy evolution, driven by the random sampling of the IMF and the specific masses of Pop III stars that are formed.
It is worth noting that the observational estimates of $t_{50}$ have very large uncertainties (e.g. Brown et al., 2014), ranging from $<100$ Myr to $\sim 1$ Gyr, and they therefore provide very limited constraints. Our predicted values are systematically short ($\sim 30$ Myr), however, the $t_{50}$ values can be adjusted by including an additional free parameter, i.e., the gas infall time, here equal to $t_{\rm inf}\approx t_{\rm ff}(z)/4\approx 15$ Myr (Rossi et al., 2021). By increasing $t_{\rm inf}$ we can obtain a better agreement with the observed $t_{50}$ without changes in the total luminosity nor the MDF (see Ciabattini et al. in prep.).
The MDF plays a key role in constraining the models, as it is highly sensitive to both the Pop III EDF and the choice of Pop II stellar yields/EDFs. In Fig.2, the observed MDF of Boötes I (Jenkins et al., 2021) is compared with the predicted MDFs for various combinations of Pop III EDFs and Pop II stellar yield sets (LC, NK CC, NK CC+HN). Each panel corresponds to a specific combination. The contribution to the total chi-squared arising specifically from the MDF is reported in the column $\chi^{2}_{\rm MDF}$ of Table 1.